# Origin of Pulsed Radio Emission from Magnetars

### v1

@@ 1 Introduction
The remarkable magnetospheric activity of magnetars likely results from their surface deformations driven by internal magnetic stresses (Kaspi & Beloborodov, 2017). These motions twist the magnetospheric field lines, creating $\nabla\times\boldsymbol{B}\parallel\boldsymbol{B}$ in the nearly force-free magnetosphere (Thompson et al., 2002; Beloborodov & Thompson, 2007). The twist of magnetic field $\boldsymbol{B}$ is sustained by electric current $\boldsymbol{j}_{0}=(c/4\pi)\nabla\times\boldsymbol{B}$, which has a long lifetime (1-10 yr) on field lines that extend far from the magnetar (Beloborodov, 2009). The gradual resistive decay of the magnetospheric twist explains the pulsating hard X-ray emission from magnetars (Beloborodov, 2013; Hascoët et al., 2014) and likely also powers their radio pulsations.
Among the 30 known magnetars, six have been observed to enter extended periods of radio pulsations. [^1] https://www.physics.mcgill.ca/~pulsar/magnetar/main.html They exhibit properties substantially different from those of ordinary radio pulsars. First, the width of observed radio pulses is significantly larger than the small angular size of the open field line bundle (e.g. Levin et al., 2010), sometimes covering a substantial fraction of the full rotational period. Second, the radio power of magnetars, $L_{r}\sim 10^{30}\,{\rm{erg/s}}$, exceeds that of ordinary radio pulsars by about 2 orders of magnitude (e.g. Camilo et al., 2006). Third, the radio spectrum significantly evolves and can be unusually hard, extending to frequencies as large as $\sim 100{\rm GHz}$ (e.g. Camilo et al., 2008; Torne et al., 2022). Fourth, the vast majority of radio emission events occur after X-ray outbursts and last up to a few years (e.g. Esposito et al., 2020).
@@
(1) The emission is powered by enhanced dissipation near the separatrix between the open and closed twisted field lines, which extends to the light cylinder $R_{\rm LC}=cP/2\pi\sim 10^{10}$ cm (Thompson, 2008), where $P=2\pi/\Omega$ is the rotation period of the magnetar. This picture invokes an electric current $I_{\rm LC}\sim\mu\Omega^{2}/c$, where $\mu\sim B_{*}R^{3}_{*}/2$ is the magnetospheric dipole moment, $B_{*}\sim 10^{15}{\rm G}$ is the magnetic field strength at the magnetar pole, and $R_{*}\sim 10\,{\rm km}$ is the stellar radius. Powering the observed radio emission by the dissipation rate
@@
with a reasonable efficiency requires a high voltage $\Phi$, far exceeding the typical $\Phi\sim 1$ GV that sustains electric currents in the inner magnetosphere (Beloborodov & Thompson, 2007). An increase of $\Phi$ toward the light cylinder would also lead to a quick resistive untwisting of the outer magnetosphere (Beloborodov, 2009), which makes it challenging to explain the long lifetime of radio emission.
(2) The emission comes from magnetic field lines closing well inside the light cylinder (Beloborodov, 2013). This proposal is based on the realization that the plasma carrying magnetospheric currents is forced into a peculiar “radiatively-locked” state, which must sustain a two-stream instability. Beloborodov (2013) conjectured that this instability generates radio emission, however did not demonstrate the generation mechanism and its efficiency.
In this Letter, we quantitatively explore the proposal of Beloborodov (2013). As we describe below, our first-principles plasma simulations demonstrate that most of the observed properties can indeed be naturally explained by coherent emission driven by a two-stream instability in a closed, twisted flux tube originating near the magnetic pole of the magnetar.
@@
Figure 1: Two-fluid model of the plasma flow in the magnetar magnetosphere. Panels (a) and (b) show Lorentz factors of positrons, $\gamma_{+}$, and electrons, $\gamma_{-}$. In this example, we have assumed that the $e^{\pm}$ outflow is injected at $r=2R_{\star}$ with a fixed multiplicity ${\cal M}=100$ (which remains constant along the outflow). The particles move along the magnetic field lines (thick dashed green and white lines) and resonantly scatter thermal photons streaming from the surface of the magnetar. Radiation exerts a drag force on both electrons and positrons. Its local strength is characterized by the value of $D_{\pm}$ calculated in the model; the contours of $D_{\pm}=1,10,100$ are indicated by the red curves. In panel (c), we show $\gamma_{\pm}$ as well as $\gamma_{\star}$ along the thick dashed green field line. We also show velocities of electrons ($v_{d,-}^{\prime}$) and positrons ($v_{d,+}^{\prime}$) in the “radiatively locked frame”. The orange and blue shaded regions have $D_{+}\geq 3$ and $D_{-}\geq 3$, respectively.
@@
where $R_{\mathrm{max}}$ is the maximum radius reached by the magnetic field line (Beloborodov, 2009). This current is sustained by electron and positron motions along the magnetic field lines. Since magnetars rotate slowly, charge neutrality, $n_{+}\approx n_{-}=n$, is a good approximation on the closed field lines (here, $n$ is the number density, and the subscripts “++” and “-” refer to positrons and electrons, respectively). Then, the existence of a non-zero current requires a difference between the velocities of the two species, $\delta v=|v_{+}-v_{-}|\neq 0$. This difference must be sustained by an electric force against the radiative drag arising from resonant inverse Compton scattering of X-ray photons. Below we summarize a two-fluid model for the current in the presence of radiative drag that was constructed in Beloborodov (2013).
@@
$n={\cal M}\frac{|j_{0}|}{e(|v_{+}|+|v_{-}|)}\approx\frac{c\psi}{4\pi e}\frac{{\cal M}B}{(|v_{+}|+|v_{-}|)R_{\mathrm{max}}}.$ (3)
@@
It is easy to find $\gamma_{\pm}$ by integrating Equation (6) along the field lines.
For example, suppose that the $e^{\pm}$ plasma is injected at radius $r=2R_{*}$ with fixed multiplicity ${\cal M}=100$ and high Lorentz factor $\gamma_{+}|_{l=0}=100$. Assuming the magnetic field strength on the pole is $B_{*}=3\times 10^{14}\,\mathrm{G}$, the computed $\gamma_{+}$ and $\gamma_{-}$ are shown in panels (a) and (b) of Figure 1. Only field lines with apex radii $R_{\rm{max}}\geq 10R_{*}$ are considered since plasma flows on field lines with small $R_{\max}$ likely have small multiplicities, ${\cal M}\sim 1$, and also tend to lose currents through the resistive untwisting process (Beloborodov, 2009). One can see that the radiative drag force becomes strong at $r\gtrsim 10R_{*}$, pushing $\gamma_{\pm}$ towards $\gamma_{\star}$. This process is further demonstrated in panel (c), where we plot $\gamma_{\pm}$ along a selected field line shown by the thick green dashed line. The strength of the drag can be quantified using the dimensionless parameter,
@@
where $p$ is the momenta of the particles, $p_{\star}\equiv\gamma_{\star}m_{e}v_{\star}$, and $\gamma_{\star}\equiv 1/\sqrt{1-\beta_{\star}^{2}}$. Parameter $D$ represents the ratio of the local light crossing time, $r/c$, to the cooling time, $t_{\rm{cool}}\equiv-(p-p_{\star})/F_{\rm{rad}}$.222 Note the difference of this definition from $D$ defined in Beloborodov (2013). The contours of $D_{\pm}$ shown in panels (a) and (b) indicate that cooling is dynamically strong, $D_{\pm}\gtrsim 1$, in the region $10\lesssim r/R_{*}\lesssim 50$. In the polar field line zone where we find bright radio emission (see Section 6), $1\lesssim D\lesssim 100$.
@@ 3 Simulation Method and Setup
As discussed in Section 2, the cooling time is much longer than the plasma period, which leads to both a small power of the coherent emission and slow growth of the cooling-induced instability. These conditions present a challenge for the standard particle-in-cell (PIC) plasma simulation technique, as these weak signals can be overwhelmed by Poisson noise. Therefore, we decided to simulate the cooling-driven instability using a relativistic extension of the continuum code Gkeyll (Juno et al., 2018; Hakim & Juno, 2020), which solves the Vlasov-Maxwell system of equations using a discontinuous Galerkin scheme (Juno et al., in preparation). Results of our simulations are numerically converged for all aspects of the analysis. Simulations are performed in the RL frame. In Section 2, quantities measured in this frame were denoted with a prime; however, in Sections 3-5 below we will omit the prime to simplify the notation, as we do not use any other frames in these sections.
@@
where $c(\nabla\times\boldsymbol{B}_{\rm tw})/4\pi=j_{0}\hat{x}=-2n_{e}ev_{d}$. One-dimensional simulations are electrostatic and do not include the field perturbation $\boldsymbol{B}$, which corresponds to setting $\boldsymbol{B}=0$ in Equation (13). Two-dimensional simulations track the evolving $\boldsymbol{B}$ using the Faraday’s induction law, $\partial\boldsymbol{B}/{\partial t}=-c\nabla\times\boldsymbol{E}$.
@@
following the procedure described in Johnson et al. (2025). Here, $\gamma_{d}\equiv 1/\sqrt{1-\beta_{d}^{2}}$, $K_{1}$ is the Macdonald function, $\xi\equiv m_{e}c^{2}/k_{B}T$, and $T$ is the plasma temperature.
Plasma in the simulation box is subject to the drag force of the form given in Equation (9). Hereafter, the radiative drag is also called “cooling,” since it tends to shrink the distribution function toward $p=0$ in the RL frame. The cooling strength is set uniform across the simulation domain. To keep the timespan of simulations reasonable, we choose the cooling time so that $\omega_{p}t_{\rm cool}=300,1000,3000$. While these values are smaller than expected in magnetars (see Eq. 11), the results of our simulations can be reliably extrapolated to realistic conditions.
@@
In 1X1V, we focus on the simulations with $v_{d}/c=0.5$, $k_{B}T/m_{e}c^{2}=0.3$, and no initial perturbation in the electric field. The extent of the configuration space is $L_{x}/2,L_{x}/2]=[-64d_{e},64d_{e}]$, where $d_{e}\equiv c/\omega_{p}=\sqrt{m_{e}c^{2}/4\pi n_{0}e^{2}}$ is the cold skin depth, and the extent of the momentum space is $64m_{e}c,64m_{e}c]$. The configuration space is resolved by $N_{x}=512$ cells, and the momentum space is resolved by $NV_{x}=1024$ cells. [^5] Gkeyll’s newest implementation, not yet tested at the time we carried out the simulations reported in this Letter, enables the use of non-uniform momentum-space grids and higher-order third-degree polynomial basis functions, thereby reducing the number of grid points needed and lowering computational costs. We employ piecewise quadratic Serendipity elements for the discontinuous Galerkin basis expansion (Arnold & Awanou, 2011). Periodic boundary conditions are applied in the configuration space, while zero-flux boundary conditions are used in the momentum space. We have also performed simulations with $v_{d}/c=0.25,0.8$, varied initial temperatures, and a strong initial electric field perturbation.
In 2X1V, our numerical setup allows the escape of waves in the $y$-direction, which is transverse to the magnetic field lines. The current-carrying zone has a finite $y$-extent. Outside this zone, we first gradually decrease $v_{d}$ and $j_{0}$ to zero (which we refer to as current transition zones), and then reduce the plasma density (forming density transition zones). The density is decreased to zero near the $y$-boundaries, so that the plasma-filled region of the computational box is separated from the boundaries by vacuum zones. All these zones are marked by the black and white horizontal lines in Figure 3. Furthermore, near the $y$-boundaries, we introduce an absorbing layer to damp escaping electromagnetic waves. This setup enables waves generated in the central current zone to propagate towards the vacuum zones and leave the plasma. We use the Poynting flux of waves escaping through the absorbing layers as a proxy for the luminosity of the low-frequency electromagnetic emission generated by turbulent plasma.
@@
This saturated state is characterized by a strong stochastic electric field, $E_{x}$ (panel a). Despite the large amplitude of $E_{x}$, its average over space and time, $\langle E_{x}\rangle$, remains comparable to $E_{0}$. Furthermore, the observed time-averaged rate of ohmic dissipation is consistent with effective $E_{x}\sim E_{0}$: $\langle\textbf{j}\cdot\textbf{E}_{x}\rangle\sim j_{0}E_{0}$. [^6] Rephrasing the same result, the dissipation due to turbulent resistivity, $\langle\delta\boldsymbol{j}\cdot\delta\boldsymbol{E}\rangle$, may be neglected compared to the dissipation driven by the average electric field $E_{0}$. Thus, the large turbulent voltage fluctuations do not affect the net dissipation rate and the corresponding lifetime of the magnetospheric twist, in agreement with Beloborodov (2013).
We also find that the ratio of the electrostatic energy density $E^{2}/8\pi$ to the plasma kinetic energy ${\cal E}_{K}=(\langle\gamma\rangle-1)n_{0}m_{e}c^{2}$ is small in the steady state, $\epsilon_{E}\equiv E^{2}/(8\pi{\cal E}_{K})\ll 1$. Nevertheless, the sustained electrostatic turbulence is not weak, as it remains marginally capable of trapping particles. This regime approximately corresponds to the so-called “critical balance” regime of electrostatic turbulence (Ewart et al., 2025; Nastac et al., 2025), where the timescale of the plasma oscillation in the electrostatic wave is comparable to the timescale for nonlinear effects. Using the corresponding terms of the kinetic equation, one can express this condition as $k\Delta v\sim eE/(m_{e}\Delta v)$, where $\Delta v$ is the saturated spread of the distribution function. Particle trapping is clearly observed in phase space, e.g., in the snapshot shown in Figure 2 (panel b). It is seen as coherent holes with a spatial scale $\sim 5d_{e}$. The appearance of coherent structures is ubiquitous in simulations for all cooling times we consider.
@@
Figure 3: Snapshot during the quasi-steady state of the 2X1V simulation for $v_{d}/c=0.5$, $\omega_{p}t_{\rm cool}=3000$. Panels (a)-(c) show the snapshots of plasma density, $n_{+}+n_{-}$, and electric field components, $E_{x}$, $E_{y}$, displaying the excitation of turbulence in the current-carrying zone, as well as the production and propagation of electromagnetic waves. In panel (a), we mark various zones of our simulation setup (see Section 3 the description) with horizontal black and white lines. The inset (d) shows density fluctuations in the current-carrying zone, and insets (e)-(f) show electromagnetic waves escaping through the absorbing layers.
@@
To study the normal modes generated by the turbulence, we have calculated the temporal and spatial Fourier transforms of the fields $E_{x}$, $E_{y}$, and $B_{z}$ that represent ordinary polarization. Their Fourier amplitudes are expected to peak on the dispersion curve $\omega(\boldsymbol{k})$ of the normal modes. Figure 4 shows the intensities of Fourier components of $E_{x}$ and $B_{z}$, $|E_{x}(\omega,{\textbf{k}})|^{2},|B_{z}(\omega,{\textbf{k}})|^{2}$, on the plane of $\omega$ and $k=\sqrt{k_{x}^{2}+k_{y}^{2}}$ for a fixed propagation angle $\theta_{k}$ (different panels correspond to $\theta_{k}=0,\pi/6,\pi/3$). For comparison, we also show the 1D simulation results. One can see that the perturbations of $E_{x}$ and $B_{z}$ do peak along the expected dispersion relation for the two well-known branches of the ordinary mode (Arons & Barnard, 1986): superluminal and subluminal (Alfvén), here computed for the cold $e^{\pm}$ beams with momenta $\pm p_{\rm peak}$. The Fourier spectrum of $E_{x}$ also shows a third branch at lower frequencies. Excitation of this mode is strongest for the field-aligned propagation, $\theta_{k}\sim 0$; it is also present in the 1X1V simulations. We associate these fluctuations with additional electrostatic modes sourced by the plateau of the distribution function, where $\partial f/\partial p_{x}\sim 0$, similar to “electron acoustic waves” (e.g., Valentini et al., 2012). [^10] In particular, we have reproduced the existence of these modes by calculating linear dispersion characteristics of 1D plasma containing two beams and a plateau (Popova, Zeng et al., in preparation).

### v2

@@ 1 Introduction
The remarkable magnetospheric activity of magnetars likely results from their surface deformations driven by internal magnetic stresses (kaspi&beloborodov_2017). These motions twist the magnetospheric field lines, creating $\nabla\times\boldsymbol{B}\parallel\boldsymbol{B}$ in the nearly force-free magnetosphere (Thompson_2002; Beloborodov_2007). The twist of magnetic field $\boldsymbol{B}$ is sustained by electric current $\boldsymbol{j}_{0}=(c/4\pi)\nabla\times\boldsymbol{B}$, which has a long lifetime (1-10 yr) on field lines that extend far from the magnetar (Beloborodov_2009). The gradual resistive decay of the magnetospheric twist explains the pulsating hard X-ray emission from magnetars (Beloborodov_2013a; Hascoet_2014) and likely also powers their radio pulsations.
Among the 30 known magnetars, six have been observed to enter extended periods of radio pulsations (Olausen_2014). [^1] https://www.physics.mcgill.ca/~pulsar/magnetar/main.html They exhibit properties substantially different from those of ordinary radio pulsars. First, the width of observed radio pulses is significantly larger than the small angular size of the open field line bundle (e.g. Levin_2010), sometimes covering a substantial fraction of the full rotational period. Second, the radio power of magnetars, $L_{r}\sim 10^{30}\,{\rm{erg/s}}$, exceeds that of ordinary radio pulsars by about 2 orders of magnitude (e.g. Camilo_2006). Third, the radio spectrum significantly evolves and can be unusually hard, extending to frequencies as large as $\sim 100{\rm GHz}$ (e.g. Camilo_2008; Torne_2022). Fourth, the vast majority of radio emission events occur after X-ray outbursts and last up to a few years (e.g. Esposito_2020).
@@
(1) The emission is powered by enhanced dissipation near the separatrix between the open and closed twisted field lines, which extends to the light cylinder $R_{\rm LC}=cP/2\pi\sim 10^{10}$ cm (Thompson_2008), where $P=2\pi/\Omega$ is the rotation period of the magnetar. This picture invokes an electric current $I_{\rm LC}\sim\mu\Omega^{2}/c$, where $\mu\sim B_{*}R^{3}_{*}/2$ is the magnetospheric dipole moment, $B_{*}\sim 10^{15}{\rm G}$ is the magnetic field strength at the magnetar pole, and $R_{*}\sim 10\,{\rm km}$ is the stellar radius. Powering the observed radio emission by the dissipation rate
@@
with a reasonable efficiency requires a high voltage $\Phi$, far exceeding the typical $\Phi\sim 1$ GV that sustains electric currents in the inner magnetosphere (Beloborodov_2007). An increase of $\Phi$ toward the light cylinder would also lead to a quick resistive untwisting of the outer magnetosphere (Beloborodov_2009), which makes it challenging to explain the long lifetime of radio emission.
(2) The emission comes from magnetic field lines closing well inside the light cylinder (Beloborodov_2013). This proposal is based on the realization that the plasma carrying magnetospheric currents is forced into a peculiar “radiatively-locked” state, which must sustain a two-stream instability. Beloborodov_2013 conjectured that this instability generates radio emission, however did not demonstrate the generation mechanism and its efficiency.
In this Letter, we quantitatively explore the proposal of Beloborodov_2013. As we describe below, our first-principles plasma simulations demonstrate that most of the observed properties can indeed be naturally explained by coherent emission driven by a two-stream instability in a closed, twisted flux tube originating near the magnetic pole of the magnetar.
@@
Figure 1: Two-fluid model of the plasma flow in the magnetar magnetosphere. Panels (a) and (b) show Lorentz factors of positrons, $\gamma_{+}$, and electrons, $\gamma_{-}$. In this example, we have assumed that the $e^{\pm}$ outflow is injected at $r=2R_{\star}$ with a fixed multiplicity ${\cal M}=100$ (which remains constant along the outflow). The particles move along the magnetic field lines (thick dashed green and white lines) and resonantly scatter thermal photons streaming from the surface of the magnetar. Radiation exerts a drag force on both electrons and positrons. Its local strength is characterized by the value of $D_{\pm}$ calculated in the model; the contours of $D_{\pm}=1,10,100$ are indicated by the red curves. In panel (c), we show $\gamma_{\pm}$ as well as $\gamma_{\star}$ along the thick dashed green field line, which has $R_{\rm max}=900R_{*}$. The evolution of $\gamma_{\pm}$, when only the radiative force is considered, is shown with dashed lines. We also show velocities of electrons ($v_{d,-}^{\prime}$) and positrons ($v_{d,+}^{\prime}$) in the “radiatively locked frame”. The orange and blue shaded regions have $D_{+}\geq 3$ and $D_{-}\geq 3$, respectively.
@@
where $R_{\mathrm{max}}$ is the maximum radius reached by the magnetic field line (Beloborodov_2009). This current is sustained by electron and positron motions along the magnetic field lines. Since magnetars rotate slowly, charge neutrality, $n_{+}\approx n_{-}=n$, is a good approximation on the closed field lines (here, $n$ is the number density, and the subscripts “++” and “-” refer to positrons and electrons, respectively). Then, the existence of a non-zero current requires a difference between the velocities of the two species, $\delta v=|v_{+}-v_{-}|\neq 0$. This difference must be sustained by an electric force against the radiative drag arising from resonant inverse Compton scattering of X-ray photons. Below we summarize a two-fluid model for the current in the presence of radiative drag that was constructed in Beloborodov_2013.
@@
$n={\cal M}\frac{|j_{0}|}{e(|v_{+}|+|v_{-}|)}\approx\frac{\psi}{4\pi e}\frac{{\cal M}B}{(|\beta_{+}|+|\beta_{-}|)R_{\mathrm{max}}},$ (3)
@@
where $m_{e}$ is the electron mass, $\gamma_{\pm}$ is the Lorentz factor of positrons/electrons and $l$ is the length along a magnetic field line. Assuming thermal X-ray emission from the magnetar surface as the only source of radiation drag, the form of $F_{\rm{rad}}(\gamma)$ is derived in Beloborodov_2013,
$F_{\rm rad}=\frac{\alpha^{2}}{4}\frac{m_{e}c^{2}}{r_{e}}\left(\frac{R_{*}}{r}\right)^{2}\left(\frac{k_{B}T_{s}}{m_{e}c^{2}}\right)^{3}\gamma g(y)(\beta_{\star}-\beta),$ (6)
where $\alpha=e^{2}/\hbar c=1/137$, $r_{e}=e^{2}/m_{e}c^{2}$ is the classical electron radius, $T_{s}$ is the surface temperature, $k_{B}$ is the Boltzmann constant, $g(y)$ is the dimensionless Planck function evaluated at the resonant frequency $\omega_{\rm res}=\gamma^{-1}(1-\beta\cos\phi)^{-1}\omega_{B}$,
$g(y)\equiv\frac{y^{3}}{e^{y}-1},\quad y\equiv\frac{\hbar\omega_{\rm res}}{k_{B}T_{s}}.$ (7)
@@
It is easy to find $\gamma_{\pm}$ by integrating Eq. (8) along the field lines.
For example, suppose that the $e^{\pm}$ plasma is injected at radius $r=2R_{*}$ with fixed multiplicity ${\cal M}=100$ and high Lorentz factor $\gamma_{+}|_{l=0}=100$. The initial Lorentz factor of electrons, $\gamma_{-}$, is determined by the relation $1-\beta_{-}/\beta_{+}=2/(\mathcal{M}+1)$, which follows directly from the definition of $\mathcal{M}$ under the assumption of charge neutrality. Assuming the magnetic field strength on the pole is $B_{*}=3\times 10^{14}\,\mathrm{G}$, the computed $\gamma_{+}$ and $\gamma_{-}$ are shown in panels (a) and (b) of Figure 1. Only field lines with apex radii $R_{\rm{max}}\geq 10R_{*}$ are considered since plasma flows on field lines with small $R_{\max}$ likely have small multiplicities, ${\cal M}\sim 1$, and also tend to lose currents through the resistive untwisting process (Beloborodov_2009). One can see that the radiative drag force becomes strong at $r\gtrsim 10R_{*}$, pushing $\gamma_{\pm}$ towards $\gamma_{\star}$. This process is further demonstrated in panel (c), where we plot $\gamma_{\pm}$ along a selected field line shown by the thick green dashed line, which has $R_{\rm max}=900R_{*}$. When only the radiative force is considered (i.e., neglecting $E_{0}$ in Eq. 5), the evolution of $\gamma_{\pm}$ is shown with dashed lines. These clearly indicate that at $l\gtrsim 20R_{*}$, without an electric field to sustain the velocity difference, the radiative force would quickly drive $\gamma_{\pm}$ toward $\gamma_{\star}$. The strength of the drag can be quantified using the dimensionless parameter,
@@
where $p$ is the momenta of the particles, $p_{\star}\equiv\gamma_{\star}m_{e}v_{\star}$, and $\gamma_{\star}\equiv 1/\sqrt{1-\beta_{\star}^{2}}$. Parameter $D$ represents the ratio of the local light crossing time, $r/c$, to the cooling time, $t_{\rm{cool}}\equiv-(p-p_{\star})/F_{\rm{rad}}$. [^2] Note that the definition of $D$ used here differs from that in Beloborodov_2013; consequently, the $D$ contours shown here are also different. The contours of $D_{\pm}$ shown in panels (a) and (b) indicate that cooling is dynamically strong, $D_{\pm}\gtrsim 1$, in the region $10\lesssim r/R_{*}\lesssim 50$. In the polar field line zone where we find bright radio emission (see Section 6), $1\lesssim D\lesssim 100$.
@@ 3 Simulation Method and Setup
As discussed in Section 2, the cooling time is much longer than the plasma period, which leads to both a small power of the coherent emission and slow growth of the cooling-induced instability. These conditions present a challenge for the standard particle-in-cell (PIC) plasma simulation technique, as these weak signals can be overwhelmed by Poisson noise. Therefore, we decided to simulate the cooling-driven instability using a relativistic extension of the continuum code Gkeyll (Juno_2018; HakimJuno_2020), which solves the Vlasov-Maxwell system of equations using a discontinuous Galerkin scheme (Juno et al., in preparation). Results of our simulations are numerically converged for all aspects of the analysis. Simulations are performed in the RL frame. In Section 2, quantities measured in this frame were denoted with a prime; however, in Sections 3-5 below we will omit the prime to simplify the notation, as we do not use any other frames in these sections.
@@
where $c(\nabla\times\boldsymbol{B}_{\rm tw})/4\pi=j_{0}\hat{x}=-2n_{e}ev_{d}$. One-dimensional simulations are electrostatic and do not include the field perturbation $\boldsymbol{B}$, which corresponds to setting $\boldsymbol{B}=0$ in Eq. (15). Two-dimensional simulations track the evolving $\boldsymbol{B}$ using the Faraday’s induction law, $\partial\boldsymbol{B}/{\partial t}=-c\nabla\times\boldsymbol{E}$.
@@
following the procedure described in Johnson_2025. Here, $\gamma_{d}\equiv 1/\sqrt{1-\beta_{d}^{2}}$, $K_{1}$ is the Macdonald function, $\xi\equiv m_{e}c^{2}/k_{B}T$, and $T$ is the plasma temperature.
Plasma in the simulation box is subject to the drag force of the form given in Eq. (11). Hereafter, the radiative drag is also called “cooling,” since it tends to shrink the distribution function toward $p=0$ in the RL frame. The cooling strength is set uniform across the simulation domain. To keep the timespan of simulations reasonable, we choose the cooling time so that $\omega_{p}t_{\rm cool}=300,1000,3000$. While these values are smaller than expected in magnetars (see Eq. 13), the results of our simulations can be reliably extrapolated to realistic conditions.
@@
In 1X1V, we focus on the simulations with $v_{d}/c=0.5$, $k_{B}T/m_{e}c^{2}=0.3$, and no initial perturbation in the electric field. The extent of the configuration space is $L_{x}/2,L_{x}/2]=[-64d_{e},64d_{e}]$, where $d_{e}\equiv c/\omega_{p}=\sqrt{m_{e}c^{2}/4\pi n_{0}e^{2}}$ is the cold skin depth, and the extent of the momentum space is $64m_{e}c,64m_{e}c]$. The configuration space is resolved by $N_{x}=512$ cells, and the momentum space is resolved by $NV_{x}=1024$ cells. [^5] Gkeyll’s newest implementation, not yet tested at the time we carried out the simulations reported in this Letter, enables the use of non-uniform momentum-space grids and higher-order third-degree polynomial basis functions, thereby reducing the number of grid points needed and lowering computational costs. We employ piecewise quadratic Serendipity elements for the discontinuous Galerkin basis expansion (Arnold_2011). Periodic boundary conditions are applied in the configuration space, while zero-flux boundary conditions are used in the momentum space. We have also performed simulations with $v_{d}/c=0.25,0.8$, varied initial temperatures, and a strong initial electric field perturbation.
In 2X1V, our numerical setup allows the escape of waves in the $y$-direction, which is transverse to the magnetic field lines. The current-carrying zone (z3) has a finite $y$-extent. Outside this zone, we first gradually decrease $v_{d}$ and $j_{0}$ to zero (which we refer to as current transition zones, z2), and then reduce the plasma density (forming density transition zones, z1). The density is decreased to zero near the $y$-boundaries, so that the plasma-filled region of the computational box is separated from the boundaries by vacuum zones (z0). Furthermore, near the $y$-boundaries, we introduce an absorbing layer to damp escaping electromagnetic waves. This setup enables waves generated in the central current zone to propagate towards the vacuum zones and leave the plasma. We use the Poynting flux of waves escaping through the absorbing layers as a proxy for the luminosity of the low-frequency electromagnetic emission generated by turbulent plasma.
@@
This saturated state is characterized by a strong stochastic electric field, $E_{x}$ (panel a). Despite the large amplitude of $E_{x}$, its average over space and time, $\langle E_{x}\rangle$, remains comparable to $E_{0}$. Furthermore, the observed time-averaged rate of ohmic dissipation is consistent with effective $E_{x}\sim E_{0}$: $\langle\textbf{j}\cdot\textbf{E}_{x}\rangle\sim j_{0}E_{0}$. [^6] Rephrasing the same result, the dissipation due to turbulent resistivity, $\langle\delta\boldsymbol{j}\cdot\delta\boldsymbol{E}\rangle$, may be neglected compared to the dissipation driven by the average electric field $E_{0}$. Thus, the large turbulent voltage fluctuations do not affect the net dissipation rate and the corresponding lifetime of the magnetospheric twist, in agreement with Beloborodov_2013.
We also find that the ratio of the electrostatic energy density $E^{2}/8\pi$ to the plasma kinetic energy ${\cal E}_{K}=(\langle\gamma\rangle-1)n_{0}m_{e}c^{2}$ is small in the steady state, $\epsilon_{E}\equiv E^{2}/(8\pi{\cal E}_{K})\ll 1$. Nevertheless, the sustained electrostatic turbulence is not weak, as it remains marginally capable of trapping particles. This regime approximately corresponds to the so-called “critical balance” regime of electrostatic turbulence (Ewart_2025; Nastac_2025), where the timescale of the plasma oscillation in the electrostatic wave is comparable to the timescale for nonlinear effects. Using the corresponding terms of the kinetic equation, one can express this condition as $k\Delta v\sim eE/(m_{e}\Delta v)$, where $\Delta v$ is the saturated spread of the distribution function. Particle trapping is clearly observed in phase space, e.g., in the snapshot shown in Figure 2 (panel b). It is seen as coherent holes with a spatial scale $\sim 5d_{e}$. The appearance of coherent structures is ubiquitous in simulations for all cooling times we consider.
@@
Figure 3: Snapshot during the quasi-steady state of the 2X1V simulation for $v_{d}/c=0.5$, $\omega_{p}t_{\rm cool}=3000$. Panels (a)-(c) show the snapshots of plasma density, $n_{+}+n_{-}$, and electric field components, $E_{x}$, $E_{y}$, displaying the excitation of turbulence in the current-carrying zone, as well as the production and propagation of electromagnetic waves. In panel (a), different zones are indicated with dashed boxes and labeled as follows: z0 is the vacuum zone, z1 is the density transition zone, z2 is the current transition zone, and z3 is the current zone. The inset (d) shows density fluctuations in the current-carrying zone, and insets (e)-(f) show electromagnetic waves escaping through the absorbing layers.
@@
To study the normal modes generated by the turbulence, we have calculated the temporal and spatial Fourier transforms of the fields $E_{x}$, $E_{y}$, and $B_{z}$ that represent ordinary polarization. Their Fourier amplitudes are expected to peak on the dispersion curve $\omega(\boldsymbol{k})$ of the normal modes. Figure 4 shows the intensities of Fourier components of $E_{x}$ and $B_{z}$, $|E_{x}(\omega,{\textbf{k}})|^{2},|B_{z}(\omega,{\textbf{k}})|^{2}$, on the plane of $\omega$ and $k=\sqrt{k_{x}^{2}+k_{y}^{2}}$ for a fixed propagation angle $\theta_{k}$ (different panels correspond to $\theta_{k}=0,\pi/6,\pi/3$). For comparison, we also show the 1D simulation results. One can see that the perturbations of $E_{x}$ and $B_{z}$ do peak along the expected dispersion relation for the two well-known branches of the ordinary mode (Arons_1986): superluminal and subluminal (Alfvén), here computed for the cold $e^{\pm}$ beams with momenta $\pm p_{\rm peak}$. The Fourier spectrum of $E_{x}$ also shows a third branch at lower frequencies. Excitation of this mode is strongest for the field-aligned propagation, $\theta_{k}\sim 0$; it is also present in the 1X1V simulations. We associate these fluctuations with additional electrostatic modes sourced by the plateau of the distribution function, where $\partial f/\partial p_{x}\sim 0$, similar to “electron acoustic waves” (e.g., Valentini_2012). [^10] In particular, we have reproduced the existence of these modes by calculating linear dispersion characteristics of 1D plasma containing two beams and a plateau (Popova, Zeng et al., in preparation).